Probability
Get insights from 304 questions on Probability, answered by students, alumni, and experts. You may also ask and answer any question you like about Probability
- 304
Questions
- 0
Discussions
- 0
Active Users
- 4
Followers
Have a question about Probability? Ask our experts
Keep it short & simple. Type complete word. Avoid abusive language.
43. The probability that a student will pass the final examination in both English and Hindi is 0.5 and the probability of passing neither is 0.1. If the probability of passing the English examination is 0.75, what is the probability of passing the Hindi examination?
43. The probability that a student will pass the final examination in both English and Hindi is 0.5 and the probability of passing neither is 0.1. If the probability of passing the English examination is 0.75, what is the probability of passing the Hindi examination?
43. Let A: student passing in Hindi
B: student passing in English
Given, P (B) = 0.75
P (A∩B) = 0.5, passing both subject
And P (A'∩B') = 0.1, i.e., passing neither subject
P (A∪B)' = 0.1
1 – P (A∪B) = 0.1
P (A∪B) = 1 – 0.1
P (A∪B) = 0.9
Hence, P (A∪B) = P (A) + P (B) – P (A∩B)
P (A) = P (A∪B) + P (A∩B) – P (
42. In an entrance test that is graded on the basis of two examinations, the probability of a randomly chosen student passing the first examination is 0.8 and the probability of passing the second examination is 0.7. The probability of passing atleast one of them is 0.95. What is the probability of passing both?
42. In an entrance test that is graded on the basis of two examinations, the probability of a randomly chosen student passing the first examination is 0.8 and the probability of passing the second examination is 0.7. The probability of passing atleast one of them is 0.95. What is the probability of passing both?
42. Let A: Student passes 1st examination
So, P (A) = 0.8
And B: Student passes 2nd examination
So, P (B) = 0.7
Also probability of passing at least one examination is P (A∪B) = 0.95
Therefore, P (A∪B) = P (A) + P (B) – P (A∩B)
0.95 = 0.8 + 0.7 – P (A∩B)
P (A∩B) = 0.8 + 0.7 – 0.95
P (A∩B) = 0.55
Hence,
41. In Class XI of a school 40% of the students study Mathematics and 30% study Biology. 10% of the class study both Mathematics and Biology. If a student is selected at random from the class, find the probability that he will be studying Mathematics or Biology.
41. In Class XI of a school 40% of the students study Mathematics and 30% study Biology. 10% of the class study both Mathematics and Biology. If a student is selected at random from the class, find the probability that he will be studying Mathematics or Biology.
41. Given that, 40% study Mathematics, 30% study Biology and 10% study both Mathematics and Biology.
Let A: Students study Mathematics.
P (A) = 40% = .
Let B: Students study Biology.
P (B) = 30% = .
So, P (A∩B) i.e. probability of student studying both Mathematics and Biology is
P
40. A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16.Determine
(i) P(not A), (ii) P(not B) and (iii) P(A or B)
40. A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16.Determine
(i) P(not A), (ii) P(not B) and (iii) P(A or B)
40. Given P (A) = 0.42
P (B) = 0.48
P (A∩B) = 0.16
(i) P (not A) = P (A') = 1 – P (A) = 1 – 0.42 = 0.58
(ii) P (not B) = P (B') = 1 – P (B) = 1 – 0.48 = 0.52
(iii) P (A or B) = P (A∪B) = P (A) + P (B) – P (A∩B)
= 0.42 + 0.48 – 0.16
= 0.74
38. If E and F are events such that P(E) = , P(F) = and P(E and F) = , find
(i) P(E or F), (ii) P(not E and not F).
38. If E and F are events such that P(E) = , P(F) = and P(E and F) = , find
(i) P(E or F), (ii) P(not E and not F).
38. Given, P (E) =
P (F) =
P (E and F) = P (E∩F) =
(i) P (E or F) = P (E∪F) = P (E) + P (F) – P (E∩F)
(ii) P (not E and not F) = P (E'∩F') = P (E∪F)' = 1 – P (E∪F)
37. Given P(A) = and P(B) = . Find P(A or B), if A and B are mutually exclusive events.
37. Given P(A) = and P(B) = . Find P(A or B), if A and B are mutually exclusive events.
37.
Given P (A) =
P (B) =
As A and B are mutually exclusive events,
P (A∩B) = 0
Hence, P (A∪B) = P (A) + P (B) – P (A∩B)
36. Fill in the blanks in following table:
P(A) P(B) P(A ∩ B) P(A ∪ B)
(i) . . .
(ii) 0.35 . . . 0.25 0.6
(iii) 0.5 0.35 . . . 0.7
36. Fill in the blanks in following table:
P(A) P(B) P(A ∩ B) P(A ∪ B)
(i) . . .
(ii) 0.35 . . . 0.25 0.6
(iii) 0.5 0.35 . . . 0.7
36. (i) Given P (A) =
P (B) =
P (A∩B) =
So, P (A∪B) = P (A) + P (B) – P (A∩B)
P (A∪B) =
(ii) Given P (A) = 0.35
P (B) =?
P (A∩B) = 0.25
P (A∪B) = 0.6
So, P (A∪B) = P (A) + P (B) – P (A∩B)
0.6 = 0.35 + P (B) – 0.25
P (B) = 0.6 – 0.35 + 0.25
P (B) = 0.5
(iii) Given P (A) = 0.5
P (B) = 0.3
35. Check whether the following probabilities P(A) and P(B) are consistently defined
(i) P(A) = 0.5, P(B) = 0.7, P(A ∩ B) = 0.6
(ii) P(A) = 0.5, P(B) = 0.4, P(A ∪ B) = 0.8
35. Check whether the following probabilities P(A) and P(B) are consistently defined
(i) P(A) = 0.5, P(B) = 0.7, P(A ∩ B) = 0.6
(ii) P(A) = 0.5, P(B) = 0.4, P(A ∪ B) = 0.8
35. Given P (A) = 0.5
P (B) = 0.7
And P (A∩B) = 0.6
As P (A∩B) > P (A) which is not possible.
The given probabilities are not consistently defined.
(ii) Given, P (A) = 0.5
P (B) = 0.4
And P (A∪B) = 0.8
So, P (A∪B) = P (A) + P (B) – P (A∩B)
0.8 = 0.5 + 0.4 – P (A∩B)
P (A∩B) = 0.5 + 0.4 – 0.8
P (A∩B) = 0.1
Hence
34. In a lottery, a person choses six different natural numbers at random from 1 to 20, and if these six numbers match with the six numbers already fixed by the lottery committee, he wins the prize. What is the probability of winning the prize in the game? [Hint order of the numbers is not important.]
34. In a lottery, a person choses six different natural numbers at random from 1 to 20, and if these six numbers match with the six numbers already fixed by the lottery committee, he wins the prize. What is the probability of winning the prize in the game? [Hint order of the numbers is not important.]
34.
. Since 6 numbers are to be choosen as fixed from a set a given 20 number, the sample space is
Let A: person wins the prize.
In order to win the prize the 6 number has to be correct i.e. all 6 of the number are to be choosen from fixed 6 numbers we have,
? P (A) = .
33. A letter is chosen at random from the word ‘ASSASSINATION’. Find the probability that letter is (i) A vowel (ii) A consonan
33. A letter is chosen at random from the word ‘ASSASSINATION’. Find the probability that letter is (i) A vowel (ii) A consonan
33. The sample space of word is
S = {A, S, A, S, I, N, A, T, I, O, N}
So, n (S) = 13.
(i) Let A: word is a vowel
A = {A, I, A, I, O}
So, n (A) = 6
? P (A) =
(ii) Let B: Word is a consonant
B = {S, N, T, N}
So, n (B) = 7
? P (B) = .